OpenStudy (anonymous):

solve the problem using substitution to get y y=4x-6

OpenStudy (anonymous):

Please somebody help me please. I'll do anything for a little help

OpenStudy (anonymous):

you need a system of equations to solve this. in other words, where is your second equation.

OpenStudy (anonymous):

the second equation is y=-4x=2

OpenStudy (ja1):

y=-4x=2 ? two = signs?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that can not be right. y=-4x=2 y = 2 -4x = 2 x = -1/2 y=4x-6 2 = 4*(-1/2) -6 2 = -2 -6 = -8 2 does not equal -8

OpenStudy (anonymous):

Wait it's y=4x-6